How many mL of 0.1566 M HCl solution would be needed to react with 2.88 grams of Ba(OH)2? Then, if a 41.83 mL beaker of 0.2980 M NaOH reacts with 25.00 mL of HNO3 solution, what is the concentration of the HNO3 solution?
Answer #1

no of moles of Ba(OH)2 = W/G.M.Wt

                                         = 2.88/171.34   = 0.0168moles

2HCl(aq) + Ba(OH)2(aq) ----------------> BaCl2(aq) + 2H2O(l)

1 mole of Ba(OH)2 react with 2 moles of HCl

0.0168 moles of Ba(OH)2 react with = 2*0.0168/1   = 0.0336moles of HCl

no of moles of HCl   = molarity* volume in L

        0.0336               = 0.1566*volume in L

volume in L    = 0.0336/0.1566   = 0.215L  

the volume of HCl   = 215ml >>>>answer

part-B

NaOH(aq) + HNO3(aq) --------------> NaNO3(aq) + H2O(l)

1 mole           1 mole

NaOH                                           HNO3

M1 = 0.2980M                             M2 =

V1 = 41.83ml                               V2 = 25ml

n1   = 1                                          n2 = 1

          M1V1/n1     =      M2V2/n2

               M2          = M1V1n2/V2n1

                               = 0.2980*41.83*1/(25*1)

                                = 0.5M

The concentration of the HNO3 solution   = 0.5M >>>>answer